OpenStudy (anonymous):

The function f(x) = x2. The graph of g(x) is f(x) translated to the left 6 units and down 5 units. What is the function rule for g(x)?

OpenStudy (anonymous):

A. g(x) = (x + 6)2 - 5 B. g(x) = (x - 5)2 + 6 C. g(x) = (x - 6)2 + 5 D. g(x) = (x + 5)2 - 6

OpenStudy (anonymous):

@satellite73 last one :)

OpenStudy (dan815):

nope

OpenStudy (anonymous):

it would be a right ?

OpenStudy (dan815):

ya

OpenStudy (anonymous):

tnx for checking my work! apprciate it

OpenStudy (anonymous):

f(x)= x2 (x-b)2+a a moves it up b moves it right.

OpenStudy (anonymous):

ANNOYING WHEN PEOPLE LIE... thank u @timo86m

OpenStudy (anonymous):

@dan815 sry u were wrong

OpenStudy (dan815):

lol

OpenStudy (anonymous):

yea soo funny . wasnt asking for a answer was checking if my anser was right no wonder y u have a high score bc u give everyone fake answers LOL @dan815

OpenStudy (dan815):

xD

OpenStudy (dan815):

what was the answer?

OpenStudy (dan815):

i actually did think about it and say A lol

OpenStudy (anonymous):

b

OpenStudy (dan815):

ah b is wrong

OpenStudy (dan815):

was it an online test?

OpenStudy (dan815):

@Jhannybean

OpenStudy (dan815):

ohh I see! you misunderstood what timo was trying to explain to you

OpenStudy (dan815):

he is talking about the a and b in the equation

OpenStudy (dan815):

your answer is really A, trust me!!

OpenStudy (jhannybean):

Translated left 6 and down 5. To get this,we take our original equation \(\large f(x)=x^2\) we go LEFT 6,that would be -6. and down 5. We can rewrite this as \[\large g(x)=(x-\color{green}{(-6)})^2 +\color{blue}{(-5)} = (x+6)^2 -5\] Our new vertex will be (-6,-5) for g(x) whereas the vertex of f(x) = (0,0)

OpenStudy (jhannybean):

http://fooplot.com/plot/qg3rb77cy2 our original and translated function :)

OpenStudy (jhannybean):

Here is your graph of answer b. Does this seem like it's moving LEFT 6 and DOWN 5 to you? http://fooplot.com/plot/eailn8ldt0

OpenStudy (dan815):

@chris00

OpenStudy (dan815):

hello jjanny have u met chris

OpenStudy (dan815):

hes my brother

OpenStudy (jhannybean):

Oh really.

OpenStudy (anonymous):

yep

OpenStudy (dan815):

we have the same croc mother

OpenStudy (anonymous):

i have to wipe his arse when he goes to the toilet...mutal agreement we made

OpenStudy (anonymous):

hahaha i just read it

OpenStudy (jhannybean):

You guys are in your own world.

OpenStudy (anonymous):

wouldn't it be A

OpenStudy (dan815):

yo do u know matlab?

OpenStudy (anonymous):

yea

OpenStudy (dan815):

how long would this assignment take

OpenStudy (anonymous):

don't know. never done it. but only two questions...

OpenStudy (dan815):

hopefully i got enough time xD i just hit the books on fourier series

OpenStudy (anonymous):

what is fourier series

OpenStudy (dan815):

you know laplace transforms and that stuff?

OpenStudy (dan815):

fourier series is pretty much the samething, except we use mostly cos and sin graphs to represent a function from what i learnt

OpenStudy (anonymous):

nup

OpenStudy (dan815):

its kindd of like a bit of linear algebra

OpenStudy (dan815):

you know how a vector can be represented by any combinations of 3 vectors that are orthogonal to each other

OpenStudy (anonymous):

yep

OpenStudy (dan815):

okay well thats the theory, suppose there are orthogonal functions infinite orthogonal functions on a given interval

OpenStudy (dan815):

than any given function can be represented as a sum of these orthogonal functions

OpenStudy (dan815):

and cosnx and sin nx where n goes from -inf to inf are such orthgonal functions

OpenStudy (dan815):

basically fourier series just represents this